OpenStudy (anonymous):

Solve the triangle. a = 12, b = 22, C = 95° c ≈ 26, A ≈ 27.6°, B ≈ 57.4° c ≈ 26, A ≈ 54.4°, B ≈ 30.6° c ≈ 25.1, A ≈ 27.6°, B ≈ 57.4° c ≈ 25.1, A ≈ 31.6°, B ≈ 53.4°

OpenStudy (anonymous):

law of cosines for this one

OpenStudy (anonymous):

\[c^2+a^2+b^2-2ab\cos(C)\] and a calculator

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

well that is because i made a mistake

OpenStudy (anonymous):

\[c^2=a^2+b^2-2ab\cos(C)\] is what i meant to write

OpenStudy (anonymous):

you need a calculator to compute \(c\)

OpenStudy (anonymous):

ok c = 25.1 that leaves me with one of these two options how do i find the A or B c ≈ 25.1, A ≈ 27.6°, B ≈ 57.4° c ≈ 25.1, A ≈ 31.6°, B ≈ 53.4°

OpenStudy (anonymous):

use the law of sines

OpenStudy (anonymous):

\[\frac{\sin(A)}{a}=\frac{\sin(C)}{c}\]

OpenStudy (anonymous):

you know three out of those four numbers, solve for \(A\)

OpenStudy (anonymous):

\[\frac{ \sin A }{ 12 } = \frac{ \sin (95) }{ 25.1 }\]

OpenStudy (anonymous):

\[\frac{ Sin A (12) }{ 12 } = \frac{ \sin 95 (12)}{ 25.1 } \] \frac{ Sin A }{ 12 } = \frac{ 11.95 }{ 25.1 } Sin A= 0.48 ??

OpenStudy (anonymous):

approximately, yes

OpenStudy (anonymous):

so \(A=\sin^{-1}(.48)\)

OpenStudy (anonymous):

28.69 ??

OpenStudy (anonymous):

i am getting 57 maybe i am doing something wrong

OpenStudy (anonymous):

for one thing, i am getting \(25.5\) for \(c\) maybe i have the numbers wrong

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

then \[\frac{ \sin A }{ 12 } = \frac{ \sin (95) }{ 26 }\]

OpenStudy (anonymous):

0.46?

OpenStudy (anonymous):

yes that is what i am getting

OpenStudy (anonymous):

then \(A=\sin^{-1}(.46)\)

OpenStudy (anonymous):

27.39

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=arcsine%28+ \frac{+12\sin+%2895%29+}{+26++}%29

OpenStudy (anonymous):

yes, go with that one

OpenStudy (anonymous):

ok thank you so much! :)